OpenStudy (anonymous):

the function y=290000(0.92)^x represents the value of an old home that has been abandoned by its owners x years ago. Find the decay rate of the old home. @Luigi0210

OpenStudy (anonymous):

is it 8%

OpenStudy (shamil98):

y = 290000(0.92)^x substitute values of x to find y if x = 1 , x = 2 , x =3 and so on. I think.

OpenStudy (shamil98):

then compare the values of the years

OpenStudy (anonymous):

yup its 8%

OpenStudy (shamil98):

okay

OpenStudy (anonymous):

do you live in united states??

OpenStudy (shamil98):

yep

OpenStudy (ranga):

If an exponential decay is written in the form: y = a(1 - r)^x then r will be the rate of decay. So y = 290000(0.92)^x can be written as y = 290000(1 - 0.08)^x r = 0.08 or 8%

OpenStudy (anonymous):

thanks sir.

OpenStudy (ranga):

you are welcome.